OpenStudy (anonymous):

what is the slope of the line that passes through (-2,-3) and (1,1)? a. 1/3 b. 1 c. 2 d.4/3

OpenStudy (mathstudent55):

Can you subtract the y-coordinates of the two points?

OpenStudy (ccswims):

Well to find slope we must do average rate of change

OpenStudy (anonymous):

\[ \frac {1-(-3)}{1 - (-2)} = \frac {4}{3} \]

OpenStudy (perl):

slope formula